# RaLiBEV: Radar and LiDAR BEV Fusion Learning for Anchor Box Free Object Detection Systems

### 01 Dec 2023 v4

@@ Abstract
In autonomous driving, LiDAR and radar play important roles in the perception of the surrounding environment. LiDAR provides accurate 3D spatial sensing information but cannot work in adverse weather like fog. On the other hand, the radar signal can be diffracted when encountering raindrops or mist particles thanks to its wavelength, but it suffers from large noise. Recent state-of-the-art works reveal that fusion of radar and LiDAR can lead to robust detection in adverse weather. The existing works adopt convolutional neural network architecture to extract features from each sensor data, then align and aggregate the two branch features to predict object detection results. However, these methods have low accuracy of bounding box estimations due to a simple design of label assignment and fusion strategies. In this paper, we propose a bird’s-eye view fusion learning-based anchor box-free object detection system, which fuses the feature derived from the radar range-azimuth heatmap and the LiDAR point cloud to estimate possible objects. Different label assignment strategies have been designed to facilitate the consistency between the classification of foreground or background anchor points and the corresponding bounding box regressions. Furthermore, the performance of the proposed object detector is further enhanced by employing a novel interactive transformer module. The superior performance of the methods proposed in this paper has been demonstrated using the recently published Oxford Radar RobotCar dataset. Our system’s average precision significantly outperforms the state-of-the-art method by 13.1% and 19.0% at IoU of 0.8 under ’Clear+Foggy’ training conditions for ’Clear’ and ’Foggy’ testing, respectively.
@@
ADAS, autonomous driving, anchor box free object detection, LiDAR, point cloud, radar, range-azimuth heatmap, label assignment, BEV fusion, interactive transformer, deep learning
@@ I Introduction
Autonomous driving systems typically employ sensors such as LiDAR, cameras, and radar to construct their perception functionality, which serves as the essential for the downstream tasks like trajectory prediction and ego motion planning [1, 2], etc. Such perception scheme has been widely demonstrated in numerous works, covering various aspects such as data collection/generation [3, 4], algorithm development [5], and real-world testing [6]. LiDAR, a key sensor used in Advanced Driver Assistance Systems (ADAS) and autonomous driving, offers rich semantic and geometric information for environmental perception tasks [7]. These tasks include object detection [9, 8], multi-object tracking [10, 11], ego localization [5, 12], and mapping [13]. To build a robust perception system for autonomous driving, recent works show that the fuse of LiDAR and camera at the raw data stream level can significantly improve the perception capability [14, 15]. However, the performance of either camera or LiDAR can severely drop in adverse weather, i.e., rain, snow, and fog [16, 17]. For instance, cameras can become blurred, the operating range of LiDAR might be significantly reduced, and the number of noise points might increase due to the impact of heavy snow or fog. As illustrated in Fig. 1(b), the number of noisy points around the ego vehicle increases while the far end LiDAR points disappear, compared to the LiDAR point cloud measured in clear weather, as shown in Fig. 1(a).
@@
  • A novel label assignment strategy, Gaussian Area-based Consistent Heatmap and intersection over union (IoU) Cost Positive Sample Assignment (GACHIPS), is proposed to combat the inconsistency between the classification of key points and the corresponding bounding box regression tasks in the anchor box free object detector.
  • A novel interactive query transformer-based fusion module, Dense Query Map-based Interactive BEV Fusion (DQMITBF), is proposed to fuse radar and LiDAR feature maps. With the introduction of dense learnable queries, the fusion module provides more flexibility to query the suitable features from both maps to fuse.
@@ III-A Anchor Box Free Object Detection
Following the idea of sensor fusion, aggregating features from each sensor under BEV representation is straightforward. Therefore, based on the YOLOv4 [65] backbone and neck, an anchor box-free radar and LiDAR BEV object detector, RaLiBEV, is proposed, as shown in Fig. 2. The corresponding LiDAR point cloud and radar range-azimuth heatmap are used as input data. Then the aggregated features are forwarded into the YOLOV4 backbone and neck for further processing. Finally, a multi-scale anchor box free detection head is employed to decode the 2D BEV rotating bounding boxes detection result.
@@
For the radar data, the strength of the range-azimuth heatmap represents the possibility of object existence under the BEV space. Where the object exists, the energy of the heatmap would be greater than that of the empty area. The radar data is converted from polar to Cartesian coordinates by applying bilinear interpolation, so the final structure shape of the data shows as $[W_{r},H_{r},C_{r}]$. For LiDAR data, since the point cloud is an unordered 3D point set, the data needs to be structured before subsequent operations. Therefore the method of PointPillars [9] is applied to form the input LiDAR data into a fixed shape $[W_{l},H_{l},C_{l}]$ under BEV.
After the pre-processing, $N$ convolution layers on each branch are applied to extract features without downsampling. Then two branch data are fused with an interactive transformer and sent into the YOLOv4 backbone and neck for further feature extraction.
A multi-scale anchor-free detection head is added at the end to decode the detection result. The first output channel of the detection head is a predicted heatmap, representing the object’s existence probability. The second to the last channel represents $\triangle x$, $\triangle y$, width, length, angle class, angle offset, and object class, respectively. $\triangle x$ and $\triangle y$ represents the predicted box center offset in x and y direction. Two channels are set to help angle regression. One is an angle classification channel, the other is an angle offset regression channel.
Following the state-of-the-art anchor-free detection paradigm, the anchor points of the feature map should be specified first, then start regression based on those locations. The 2D ground-truth label boxes generate a Gaussian distribution map as a supervision signal under each output head scale to train the network for heatmap estimation in a supervised manner. This method can be described by:
$G(X|\mu,\Sigma)=\frac{1}{2\pi^{\frac{d}{2}}\left|\Sigma\right|^{\frac{1}{2}}}e^{-\frac{1}{2}(X-\mu)^{T}\Sigma^{-1}(X-\mu)},$ (1)
where the $X$ is the 2D BEV spatial vector, $d$ is the dimension of $X$. $\mu$ is the mean of $X$, which is the center coordinates of ground-truth boxes. $\Sigma$ is a covariance matrix of the distribution, which can be calculated by:
$\Sigma=E(X-E(X))(X-E(X))^{T},$ (2)
where the four corners’ 2D coordinates of a box are used as $x$, and the $E(x)$ equals the center coordinates of the box. Iteration is done over all the ground-truth objects in one frame to generate the ground-truth Gaussian heatmap.
@@ III-B Label Assignment Strategies
where the $\mathcal{L}_{cls}$ is the foreground-background classification loss implemented with focal loss [66], the $\mathcal{L}_{box}$ is the box regression loss implemented with smooth $L1$ loss. The focal loss is modified with Gaussian mixture distribution weight $\mathcal{G}$:
@@
The $i,j$ range from 0 to feature map number of grid $M$, $\textbf{G}_{ij}\in[0,1]$ specifies the supervision Gaussian mixture distribution heatmap generated from Eq. (1), $p_{ij}\in[0,1]$ is the model’s estimated heatmap value for each anchor point in a feature map, and $\gamma$ is a static parameter that balances the weight of positive and negative samples.
@@
$\alpha$ is binary mask matrix that controls the box regression loss in which each element $\alpha_{i,j}$ is either 1 or 0. Typically, the binary mask matrix $\alpha$ is utilized to control which anchor points on the feature map are selected for calculating the box loss, denoting by $\alpha_{i,j}=1$. The chosen samples are referred to as positive samples, while the unselected ones are termed negative samples.
Because the box regression loss could only occur on the positive samples, while the foreground-background classification loss is computed upon every anchor point on the feature map. This leads to the problem that when decoding the predicted object, an anchor point with a higher foreground-background classification score does not always index the best bounding box regression result. This is known as the inconsistency problem in label assignments. A label assignment strategy with lesser inconsistency between foreground-background classification and box regression can decrease the model’s miss-detection rate and help achieve higher detection accuracy.
Typically, three strategies of label assignment could be applied, including bipartite matching, multi-positive label assignment, and single-positive label assignment. The ’multi/single’ indicates how many anchor points can be chosen as the positive sample for one object.
@@ III-B1 Bipartite Matching Label Assignment
The typical bipartite matching label assignment is based on the Hungarian algorithm. This set-based label assignment builds a cost matrix of IoU between predicted and ground-truth boxes, where the predicted boxes are decoded from a predefined location. Then it finds the best matching positive samples by the Hungarian algorithm. The traditional bipartite matching process is based on anchors [67], while DETR [68] leveraging object query changes this matching art to a total end-to-end style. Using the Hungarian algorithm, DETR computes loss directly between predefined object queries and ground-truth object boxes. Then after training, the converged queries can be directly decoded into the final predicted box without Non-Maximum Suppression (NMS). However, the size of the query map in the proposed method is set to be the same as the size of the radar and LiDAR feature map. Such a large number of queries leads to massive computational consumption, which is not feasible when applying the Hungarian algorithm. Thus, such bipartite matching label assignment is not considered for our proposed method.
Figure 3: Label assignment strategies. The ground-truth heatmap defines the orange-colored region of target objects. The red ellipses are prediction heatmap Gaussian distribution. The green and red dots represent the classification and regression key points.
@@ III-B2 Multi-Positives Label Assignment
Following the idea that multiple positive anchor points may alleviate positive and negative sample imbalance problems in anchor-free detection architecture, a Gaussian area-based multi-positive label assignment is designed. Based on the ground-truth Gaussian mixture distribution heatmap G created in Section III-A, a static threshold is used to extract the specific region as the positive anchor points. foreground-background classification and box regression are then applied together on each positive anchors points. As Fig. 3(b) shows, all these dots with green and red colors overlapped are the selected positive samples. They are all located in the ground-truth Gaussian distribution orange-colored area, thus this approach is named as Gaussian Area-based Multi-Positives Sample assignment.
@@ III-B3 Single Positive Label Assignment
In order to prevent numerical regression conflicts caused by multiple positive samples, single positive sample allocation strategy has been designed. The fundamental characteristic of a single positive sample is that, for each ground-truth object, there is only one positive anchor point corresponding to it, which contributes to the gradient of the loss in each training iteration. Four methods have been proposed successively to alleviate the inconsistency between foreground-background classification and box regression mentioned in the previous discussion. The experiment shows that abandoning the center prior can most effectively alleviate the consistency issue between object detection and regression.
The proposed four methods are Direct Index based Positive Sample (DIPS) assignment, Gaussian Area-based Heatmap cost Positive Sample (GAHPS) assignment, Gaussian Area-based Heatmap and IoU cost Positive Sample (GAHIPS) assignment and Gaussian Area-based Consistent Heatmap and IoU cost Positive Sample (GACHIPS) assignment. They share the same training framework as below:
@@
The first step is generating candidate anchor points. This step is the common procedure for all four label assignment strategies. According to Eq. (1), a Gaussian distribution $\bm{G}_{k},k=1,...,n$} is generated for each ground-truth object, and the multiple objects form a Gaussian mixture distribution $\mathcal{G}$. By setting a fixed threshold, the anchor points inside each Gaussian distribution can be obtained, which are called candidate anchor points. This steps can be described as:
@@
  • The second approach is termed as the Gaussian Area-based Heatmap cost Positive Sample (GAHPS) assignment. In this method, we substitute the cost calculation from determining the centers of ground-truth labels to assessing the maximum predicted heatmap value for each anchor point. The remaining steps, including the first and the fourth, are kept unchanged. Firstly, the model inferences predicted heatmap $\bm{P}^{cls}_{k}$ and boxes $\bm{P}^{box}_{k}$ follows Eq. (9). As illustrated in Fig. 3(d), the red fading ellipses in Fig. 3(d) represents the predicted Gaussian distribution heatmap. This step is described as:
@@
  • The third approach is termed as the Gaussian Area-based Heatmap and IoU cost Positive Sample (GAHIPS) assignment. In this method, we go beyond merely substituting the cost with the predicted heatmap value. Instead, we replace it with the sum of the predicted heatmap value and the IoU value between the predicted box and the ground-truth box. Therefore, the first step is to calculate the IoU score of predicted bounding box and ground-truth box.
    the decoded $\bm{P}^{box}_{k}$ derived from Eq. (9) is utilized to compute the IoU scores $\bm{S}^{IoU}_{k}$ in relation to the ground-truth bounding boxes $\bm{\mathcal{G}}^{box}_{k}$ as depicted in:
@@
    To be noted that, although the processes of determining $\bm{\alpha}$ are different in DIPS, GAHPS and GAHIPS, the ways of computing final loss are same. However, because the foreground-background classification loss and box regression loss are independent of each other, such definition of final loss leads to a consistency gap between the two loss above. To address this issue, the fourth design is proposed.
    The fourth design is Gaussian Area-based Consistent Heatmap and IoU cost Positive Sample (GACHIPS) assignment, which is a divergence of the GAHIPS. GACHIPS follows the previous three steps as the GAHIPS, but modifies the final step of loss computation. All the previous designs conceal a fact that the positive sample position of foreground-background classification is designed to converge to the center of the ground-truth box. Meanwhile, the box regression positive sample position moves with the it when training iteration by iteration. Those strategies build a weak relation between the two loss part. However, the position of positive samples is not necessarily required to be at the center of the ground-truth bounding box. Therefore, the foreground-background classification loss should performs after the positive sample positions are set together with box regression loss. This means though the other steps are exact same as GAHIPS, the former foreground-background classification loss function which is shown in Eq. (4), should be replaced by:
@@
In the end, a toy visualization of aforementioned label assignment strategies is shown in Fig. 3.
@@
Figure 4: Interactive transformer BEV fusion approaches.

### 06 Feb 2024 v5

@@ Abstract
In autonomous driving, LiDAR and radar are crucial for environmental perception. LiDAR offers precise 3D spatial sensing information but struggles in adverse weather like fog. Conversely, radar signals can penetrate rain or mist due to their specific wavelength but are prone to noise disturbances. Recent state-of-the-art works reveal that the fusion of radar and LiDAR can lead to robust detection in adverse weather. The existing works adopt convolutional neural network architecture to extract features from each sensor data, then align and aggregate the two branch features to predict object detection results. However, these methods have low accuracy of predicted bounding boxes due to a simple design of label assignment and fusion strategies. In this paper, we propose a bird’s-eye view fusion learning-based anchor box-free object detection system, which fuses the feature derived from the radar range-azimuth heatmap and the LiDAR point cloud to estimate possible objects. Different label assignment strategies have been designed to facilitate the consistency between the classification of foreground or background anchor points and the corresponding bounding box regressions. Furthermore, the performance of the proposed object detector is further enhanced by employing a novel interactive transformer module. The superior performance of the methods proposed in this paper has been demonstrated using the recently published Oxford Radar RobotCar dataset. Our system’s average precision significantly outperforms the state-of-the-art method by 13.1% and 19.0% at Intersection of Union (IoU) of 0.8 under ’Clear+Foggy’ training conditions for ’Clear’ and ’Foggy’ testing, respectively.
@@
Autonomous driving, anchor box free object detection, LiDAR, point cloud, radar, range-azimuth heatmap, label assignment, bird eye’s view fusion, interactive transformer, deep learning.
@@ I Introduction
Autonomous driving systems typically employ sensors such as LiDAR, cameras, and radar to construct their perception functionality, which serves as the essential for the downstream tasks like trajectory prediction and ego motion planning [1, 2], etc. Such perception scheme has been widely demonstrated in numerous works, covering various aspects such as data collection/generation [3, 4], algorithm development [5], and real-world testing [6]. LiDAR, a key sensor used in Advanced Driver Assistance Systems (ADAS) and autonomous driving, offers rich semantic and geometric information for environmental perception tasks [7]. These tasks include object detection [9, 8], multi-object tracking [10, 11], ego localization [5, 12], and mapping [13]. Recent works show that fusing LiDAR and camera data at the raw stream level can significantly improve the robustness of autonomous driving perception systems [14, 15]. However, the performance of either camera or LiDAR can severely drop in adverse weather, i.e., rain, snow, and fog [16, 17]. For instance, cameras can become blurred, the operating range of LiDAR might be significantly reduced, and the number of noise points might increase due to the impact of heavy snow or fog. As illustrated in Fig. 1(b), the number of noisy points around the ego vehicle increases while the far end LiDAR points disappear, compared to the LiDAR point cloud measured in clear weather, as shown in Fig. 1(a).
@@
  • A novel label assignment strategy, Gaussian Area-based Consistent Heatmap and IoU Cost Positive Sample Assignment (GACHIPS), is proposed to combat the inconsistency between the classification of key points and the corresponding bounding box regression tasks in the anchor box free object detector.
  • A novel interactive transformer-based fusion module, Dense Query Map-based Interactive BEV Fusion (DQMITBF), is proposed to fuse radar and LiDAR feature maps. With the introduction of dense, learnable queries and a symmetric fusion architecture, the fusion module provides an information-sensitive fusion mechanism to aggregate suitable features from both maps under various weather conditions.
@@ III-A Anchor Box Free Object Detection
Following the idea of sensor fusion, aggregating features from each sensor under BEV representation is straightforward. As depicted in Fig. 2, the raw LiDAR point cloud and radar range azimuth heatmap undergo pre-processing via their respective sensor branch’s feature extractors. These two features are then fed into fusion module for feature fusion. The combined features are subsequently processed by a YOLOV4-based [65] feature extractor and a multi-scale anchor-free object detection head to decode the 2D BEV rotating bounding boxes detection result.
@@
Specifically, the radar data is converted from polar to Cartesian coordinates by applying bilinear interpolation, so the final structure shape of the data shows as $[W_{r},H_{r},C_{r}]$. The $W,H,C$ denote width, height and channel of the feature map respectively. For LiDAR data, since the point cloud is an unordered 3D points set, the data needs to be structured before subsequent operations. Therefore the method of PointPillars [9] is applied to form the input LiDAR data into a fixed shape $[W_{l},H_{l},C_{l}]$ under BEV.
After the pre-processing, $N$ convolution layers on each branch are applied to extract features without downsampling. Then two branch data are fused with an interactive transformer module and sent into the YOLOv4 backbone for further feature extraction.
A multi-scale anchor-free detection head is added at the end to decode the detection results. The first output channel of the detection head is a predicted heatmap, representing the object’s existence probability. The second to the last channel represent $\triangle x$, $\triangle y$, width, length, angle classification, angle offset, and object class, respectively. $\triangle x$ and $\triangle y$ denote the predicted box center offset in x and y direction. Two channels are set to help angle regression. One is an angle classification channel, the other is an angle offset regression channel.
Following the state-of-the-art anchor-free detection paradigm, the positive anchor points of the feature map should be specified first, then start regression based on those locations. For each ground-truth bounding box in the BEV, a Gaussian distribution map is generated under each head scale. This serves as a supervision signal to train the network for heatmap prediction in a supervised learning approach. The generation method can be described by:
$\bm{G}(\bm{x}|\bm{\mu},\bm{\Sigma})=\frac{1}{2\pi^{\frac{d}{2}}\left|\bm{\Sigma}\right|^{\frac{1}{2}}}e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})},$ (1)
where the $\bm{x}=[i,j]$ is the BEV spatial vector, and $i,j$ range from 0 to number of grid on the feature map. $d$ is the dimension of $\bm{x}$. $\bm{\mu}$ is the mean of $\bm{x}$, which is the center coordinates of ground-truth boxes. $\bm{\Sigma}$ is a covariance matrix of the distribution, which can be calculated by:
$\bm{\Sigma}=E(\bm{x}-E(\bm{x}))(\bm{x}-E(\bm{x}))^{T},$ (2)
where the four corners’ 2D coordinates of a box are used as $\bm{x}$, and the $E(\bm{x})$ equals the center coordinates of the box. Such operation is done iteratively over all ground-truth objects in one time frame, thus a ground-truth Gaussian mixture distribution is generated.
@@ III-B Label Assignment Strategies
where the $\mathcal{L}_{cls}$ is the foreground-background classification loss implemented with focal loss [66], the $\mathcal{L}_{box}$ is the box regression loss implemented with smooth $L1$ loss. The focal loss is modified with Gaussian mixture distribution weight $G_{i,j}$:
@@
$M$ is the grid number of feature map. $G_{i,j}\in[0,1]$ specifies each element in the ground-truth Gaussian mixture distribution generated from Eq. (1). $p_{i,j}\in[0,1]$ is the model’s estimated value for each anchor point on BEV feature map, and $\gamma$ is a static parameter that balances the weight of positive and negative samples. By weighting the gradient with a Gaussian weight, the model’s convergence will be more stable and quicker [58].
@@
where $\alpha_{i,j}\in\{0,1\}$ is the binary element of mask matrix $\bm{\alpha}$. Typically, $\alpha$ is utilized to control which anchor points on the BEV feature map are selected for calculating the box loss. The chosen samples are referred as positive samples, denoting by $\alpha_{i,j}=1$, while the unselected ones are termed as negative samples with $\alpha_{i,j}=0$.
The box regression loss could only occur on the positive samples, while the foreground-background classification loss is computed upon every anchor point on the feature map. This leads to the problem that an anchor point with a higher foreground-background classification score does not always index the best bounding box regression result during decoding the predicted object. This is known as the inconsistency problem in label assignments. A label assignment strategy with lesser inconsistency between foreground-background classification and box regression can decrease the model’s miss-detection rate and help achieve higher detection accuracy.
Typically, three strategies of label assignment could be applied, including bipartite matching, multi-positive label assignment, and single-positive label assignment. “Multi/single” indicates how many anchor points can be chosen as the positive sample for one object.
@@ III-B1 Bipartite Matching Label Assignment
The typical bipartite matching label assignment is based on the Hungarian algorithm. This set-based label assignment builds a cost matrix of IoU between predicted and ground-truth boxes, where the predicted boxes are decoded from a predefined location. Then it finds the best matching positive samples by the Hungarian algorithm. The traditional bipartite matching process is based on anchors [67], while DETR [68] leveraging object query changes this matching art to a total end-to-end style. Using the Hungarian algorithm, DETR computes loss directly between predefined object queries and ground-truth object boxes. Then after training, the converged queries can be directly decoded into the final predicted boxes without Non-Maximum Suppression (NMS). However, the size of the query map in our proposed method is set to be the same as the size of the radar and LiDAR feature map. Such a large number of queries leads to massive computational consumption, which is not feasible when applying the Hungarian algorithm. Thus, such bipartite matching label assignment is not considered for our proposed method.
Figure 3: Overview of label assignment strategies for object detection. (a) Identifies the object with a red-bordered ground-truth bounding box and a green headline. Yellow and red ellipses represent ground-truth and predicted Gaussian distributions, respectively. Red dots are positive sample points, and green dots mark the ground-truth Gaussian centers. Strategies (b) to (e) show different methods for selecting positive samples for box loss calculation, ranging from (b) using all anchor points within the ground-truth area, to (c) selecting the center, (d) the point with the highest foreground-background classification score, and (e) the point with the highest ”score plus IoU”. Strategy (f) integrates the approach from (e) with an alternative loss function as Eq. (14).
@@ III-B2 Multi-Positives Label Assignment
Following the idea that multiple positive anchor points may alleviate positive and negative sample imbalance problems in anchor-free detection architecture, a Gaussian area-based multi-positive label assignment is designed. Based on the ground-truth Gaussian mixture distribution heatmap G created in Section III-A, a static threshold is used to extract the specific region as the positive anchor points. Foreground-background classification and box regression are then applied together on each positive anchors points. As Fig. 3(b) shows, all dots with red colors are the selected positive samples. They are all located in the ground-truth Gaussian distribution yellow-colored area, thus this approach is named as Gaussian Area-based Multi-Positives Sample assignment.
@@ III-B3 Single Positive Label Assignment
In order to prevent numerical regression conflicts caused by multiple positive samples, single positive sample allocation strategy has been designed. The fundamental characteristic of a single positive sample is that, for each ground-truth object, there is only one positive anchor point corresponding to it, which contributes to the gradient of the loss in each training iteration. Four methods have been proposed successively to alleviate the inconsistency between foreground-background classification and box regression mentioned in the previous discussion. The experiment shows that abandoning the center prior can most effectively alleviate the consistency issue between object detection and regression. The proposed four methods share the same training framework as below:
@@
The first step is generating candidate anchor points. This step is the common procedure for all four label assignment strategies. According to Eq. (1), a Gaussian distribution $\bm{G}_{k},k=1,...,n$} is generated for each ground-truth object, and the multiple objects form a Gaussian mixture distribution $\bm{G}$. By setting a fixed threshold, the anchor points inside each Gaussian distribution can be obtained, which are called candidate anchor points. This steps can be described as:
@@
  • The second approach is termed as the Gaussian Area-based Heatmap cost Positive Sample (GAHPS) assignment. In this method, we substitute the cost calculation from determining the centers of ground-truth boxes to assessing the maximum predicted heatmap value for each anchor point. The remaining steps, including the first and the fourth, are kept unchanged. Firstly, the model inferences predicted heatmap $\bm{P}^{cls}_{k}$ and boxes $\bm{P}^{box}_{k}$ follows Eq. (9). As illustrated in Fig. 3(d), the red fading ellipse represents the predicted Gaussian distribution heatmap. This step is described as:
@@
  • The third approach is termed as the Gaussian Area-based Heatmap and IoU cost Positive Sample (GAHIPS) assignment, as shown in Fig. 3(e). In this method, we go beyond merely substituting the cost with the predicted heatmap value. Instead, we replace it with the sum of the predicted heatmap value and the IoU value between the predicted box and the ground-truth box. Therefore, the first step is to calculate the IoU score of predicted bounding box and ground-truth box. The decoded $\bm{P}^{box}_{k}$ derived from Eq. (9) is utilized to compute the IoU scores $\bm{S}^{IoU}_{k}$ in relation to the ground-truth bounding boxes $\bm{\mathcal{G}}^{box}_{k}$:
@@
    To be noted that, although the processes of determining $\bm{\alpha}$ are different in DIPS, GAHPS and GAHIPS, the ways of computing final loss are same. However, because the foreground-background classification loss and box regression loss are independent from each other, such definition of final loss leads to a consistency gap between the two loss above. To address this issue, the fourth design is proposed.
  • The fourth design is Gaussian Area-based Consistent Heatmap and IoU cost Positive Sample (GACHIPS) assignment, which is a divergence of the GAHIPS. GACHIPS follows the previous three steps as the GAHIPS, but modifies the final step of loss computation as indicated in Fig. 3(f). By doing it, the foreground-background classification loss and box regression loss are both controlled by $\bm{\alpha}$. Specifically, the former foreground-background classification loss function which is shown in Eq. (4), should be replaced by:
@@
Figure 4: Comparison of direct transformer for BEV fusion with dense query map-based interactive transformer for BEV fusion.